Find the area of a rectangle if three of its corners are (-0.1,-3.4), (-7.9,-3.4), and (-0.1,6.8).
Answer by checkley71(8403)

x=-.1-(-7.9) or x=7.8 & y=-3.4-(+6.8) or y=-10.2 thus the dimentions are 7.8 by 10.2 or the area=7.9*10.2 or area=79.56